OpenStudy (anonymous):

I am a freshman in College and I am in Intermediate Algebra and I am trying to solve rational exponents and radical expressions. Can you please help me???

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

this is the problem I am working on maybe you can help me with it, 27 1/3 needs to be simplified

OpenStudy (anonymous):

\[27^{1/3}\]

OpenStudy (anonymous):

I guess the answer the teacher gave us was first you simplify it to \[\sqrt[3\sqrt{27}]{?}\] and the end result answer was 3, How does that work???

OpenStudy (anonymous):

Okay, so if I have an exponent that is a fraction, the top number works like a regular power, but the bottom number becomes like a ROOT. So, in general, I have: \[x^{a/b} = \sqrt[b]{x^{a}}\]

OpenStudy (anonymous):

yea that's the way it was explained I guess, but the answer is 3, and so I am trying to figure out how to solve these rational and radical expressions

OpenStudy (anonymous):

Well, do you see how 27^(1/3) gives me \[\sqrt[3]{27}\]

OpenStudy (anonymous):

The 1 in the top is my power, but a power of 1 doesn't do anything special. The 3 in the bottom gives me my root.

OpenStudy (precal):